OpenStudy (anonymous):

Find the quotient (5/43) / (25/86)

OpenStudy (anonymous):

@Kainui same thing on this one ?

OpenStudy (anonymous):

\[\frac{ 5 }{ 43 } \times \frac{ 86 }{ 25 }\]

OpenStudy (anonymous):

Yes, @Kainui is correct, I assume he taught you how to do divide fractions.

OpenStudy (anonymous):

yeah lol so it's (430/1075) ?

OpenStudy (anonymous):

.4 , or (2/5) is the answer